May you help me with this problem:
2p-q=2
2p+3q=22
The instructions are to use linear combinations to solve the linear system.
Then check your solution. Thank you for time.

Answer by Alan3354(69443)

Do it like this one.
------------
3j+5k=19
j-2k=-1
------------
Multiply the 2nd eqn by 3 to make the coefficients equal
3j+5k=19
3j-6k=-3
---------- Subtract
-k = 22
k = 22
-------
j - 2k = -1
j -44 = -1
j = 43